Grade 9General Physics

A body is dropped from a 100 m high cliff and at the same time another body is thrown from the ground with25 m/s velocity in upward direction.Where the two will meet

Mutayyib shakeel
8 Years agoGrade 9

1 Answer

Khimraj
8 Years ago
Let they meet at height h from ground.
Then 
100-h =  (½)gt2        – (1)
and
h = 25t  – (½)gt2      – (2)
Adding (1) and (2)
100 = 25t
then
h = 25*4 – (1/2)*10*42
h = 100 – 80 = 20m.
Hope it clears. If you like answer then please approve it.